# The Hilbert-space structure of free fermions in disguise

### v1

@@ I Introduction
Given the long history of the JW transformation, it is perhaps surprising that only recently new types of spin chains that can be mapped to free fermions were discovered. In 2019, Paul Fendley constructed a model which is quartic in the JW fermions, and still quadratic in terms of suitably defined fermionic creation and annihilation operators [19] (see Refs. [20, 21, 22] for earlier work in this direction). Remarkably, it was proven that the spin chain found by Fendley cannot be diagonalized by any generalized JW transformation [23], making the model fundamentally different from the previously known ones. This model and later generalizations [24, 25, 26, 27] are now identified with the name of free fermions in disguise (FFD).
The mapping to free fermions in FFD models gives access to the full Hamiltonian spectrum. Still, the computation of spin correlation functions is generally hard, because the dictionary between states in the spin and free-fermion Hilbert spaces is highly non-trivial. Indeed, on the one hand the FFD mapping is non-linear and non-local; on the other hand, each Hamiltonian eigenspace is known to be exponentially degenerate [19], so that fermionic Fock states cannot be uniquely associated with spin states. As a consequence, the problem of computing correlation functions in FFD models remains essentially open, with only a few results available [28]. In some respect, the current situation is similar to that of interacting integrable models [29, 30]: while the Bethe ansatz gives direct access to the Hamiltonian spectrum, computing correlation functions in arbitrary eigenstates is hard, and analytic results exist only in a few special cases [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43].
@@
This work is organized as follows. We begin in Sec. II by recalling the model introduced by Paul Fendley (Sec. II.1) and its exact solution in terms of free fermions (Sec. II.2). Our results are contained in Secs. II.3, IV, and V, and summarized in Sec III. In Sec. II.3 we show that the Hilbert space factorizes as $\mathcal{H}=\mathcal{H}_{F}\otimes\mathcal{H}_{D}$, where $\mathcal{H}_{F}$ supports the fermionic operators. The structure of $\mathcal{H}_{D}$ is further resolved in the later sections, where we show that $\mathcal{H}_{D}=\mathcal{H}_{F^{\prime}}\otimes\mathcal{H}_{\widetilde{D}}$ and characterize the two factors individually. In particular, in Sec. IV we show that $\mathcal{H}_{\widetilde{D}}$ is a “representation-dependent” subspace, generated by the common eigenstates of an extensive set of commuting Pauli strings, while in Sec. V we show that the space of operators supported on $\mathcal{H}_{F^{\prime}}$ is generated by an operator algebra of ancillary free fermion in disguise. Finally, Sec. VI contains our conclusions and an outlook for future work.
II The Model and its Hilbert-space structure
@@ II.2 Exact solution
As an important property, transfer matrices at different spectral parameters commute, $[T_{M}(u),T_{M}(u^{\prime})]=0$, and the Hamiltonian can obtained as the logarithmic derivative
@@
so that $T_{M}^{-1}(u)\propto T_{M}(-u)$. Here $P_{M}(u^{2})$ is a polynomial of degree $S=[(M+2)/3]$ in $u^{2}$ ([]delimited-[][$\cdot$][ ⋅ ] denotes the integer part), whose zeroes $\{u_{k}^{2}\}_{k=1,\ldots S}$ uniquely determine the energy spectrum of $H$: denoting $\epsilon_{k}=1/u_{k}$, there are $2^{S}$ distinct energy eigenvalues of the form
@@
$\Psi_{k}\propto T_{M}(\epsilon_{k})\leavevmode\nobreak\ \chi\leavevmode\nobreak\ T_{M}(-\epsilon_{k})\,.$ (7)
@@
II.3 The Hilbert-space structure
In this section, we use the fermionic operators (7) to show that the FFD Hilbert space can be decomposed as $\mathcal{H}=\mathcal{H}_{F}\otimes\mathcal{H}_{D}$. Here, $\mathcal{H}_{F}$ is a space of dimension $2^{S}$ that the fermionic operators are supported on, while $\mathcal{H}_{D}$ accounts for the exponentially large degeneracies observed in the spectrum of $H$. This simple observation is our first result and the basis for the subsequent analysis presented in the rest of this work, cf. Sec. III
@@ II.3 The Hilbert-space structure
$\mathcal{K}=\{\ket{v}\in\mathcal{H}\ :\ \Psi_{k}\ket{v}=0\ \leavevmode\nobreak\ \forall k=1\,,\ldots S\}\,.$ (10)
Clearly, $\mathcal{K}$ coincides with the degenerate space associated with the ground-state energy, and has dimension ${\rm dim}(\mathcal{K})=2^{M-S}$, cf Sec. II.2. We denote by {|$\{\ket{w_{j}}\}_{j=1}^{2^{M-S}}$ an orthogonal basis of $\mathcal{K}$. Next, we define
|$\ket{f_{j,\alpha}}=(\Psi^{\dagger}_{1})^{\alpha_{1}}(\Psi^{\dagger}_{2})^{\alpha_{2}}\cdots(\Psi^{\dagger}_{S})^{\alpha_{S}}\ket{w_{j}}\,,$ (11)
where $\alpha_{j}=0,1$, and $(\alpha_{1}\alpha_{2},\ldots\alpha_{S})$ is the binary representation of the integer $\alpha\in[0,\ldots 2^{S}-1]$. Because of the properties of the fermionic operators, it is immediate to see that |$\ket{f_{j,\alpha}}$ are mutually orthogonal, and thus form a basis for the Hilbert space $\mathcal{H}$.
@@
|$\ket{\alpha}=(c^{\dagger}_{1})^{\alpha_{1}}(c^{\dagger}_{2})^{\alpha_{2}}\cdots(c^{\dagger}_{S})^{\alpha_{S}}|\Omega\rangle\,,$ (13)
where $\{c_{j}^{\dagger}\}$ are canonical fermionic creation operators, |$\ket{\Omega}$ is the corresponding vacuum state, while $\mathcal{H}_{D}$ is generated by the basis vectors {|$\{\ket{j}\}_{j=1}^{2^{M-S}}$. Note that the order of the operators $\Psi_{j}$ and $c_{j}$ in the definitions (11) and (13) is fixed. It is then easy to show that $\varphi$ is an isomorphism preserving the Hilbert-Schmidt scalar product.
Next, we claim that the isomorphism $\varphi$ maps the operators $\Psi_{k}$ into operators supported on $\mathcal{H}_{F}$, namely $𝟙\varphi(\Psi_{k})=c_{k}\otimes\openone$. Indeed, by definition of $\varphi$,
@@
so that all fermionic operators are supported on $\mathcal{H}_{F}$. Finally, since the Hamiltonian is written explicitly in terms of the fermionic operators, it follows that (with a slight abuse of notation) $𝟙H=H\otimes\openone$. Therefore, $\mathcal{H}_{D}$ accounts for the exponential degeneracy of each Hamiltonian eigenstate.
III Overview of our results
In the previous section, we have shown that the Hilbert space admits the factorization $\mathcal{H}=\mathcal{H}_{F}\otimes\mathcal{H}_{D}$. In the rest of this work, we will further refine the structure of $\mathcal{H}_{D}$. Before proceeding, it is useful to provide an overview of our main results, which we report in this section.
@@ III Overview of our results
  • We will show that ${\mathcal{H}}_{F^{\prime}}$ supports the action of all operators written in terms of the generators $h_{m}$ that commute with the fermion creation/annihilation operators. We will denote the algebra of such operators by $\mathcal{A}_{F^{\prime}}$ and we will refer to it as the ancillary fermionic algebra (to be distinguished by the fermionic algebra $\mathcal{A}_{F}$ supported on $\mathcal{H}_{F})$. Importantly, the definition of $\mathcal{A}_{F^{\prime}}$ is independent of the specific representation chosen, because it can be stated entirely in terms of the generators $h_{m}$ and their algebraic relations. However, we will show that $\mathcal{A}_{F^{\prime}}$ does depend on the Hamiltonian couplings $\{b_{m}\}$. Going further, we will show in Sec. V, that $\mathcal{A}_{F^{\prime}}$ is a fermionic algebra, generated by a new family of ancillary fermion operators
@@
The dimensions of the subspaces $\mathcal{H}_{F}$, $\mathcal{H}_{F^{\prime}}$, and $\mathcal{H}_{\widetilde{D}}$ depend, respectively, on the integers $S$, $S^{\prime}$, $N_{P}$ and $N_{C}$ introduced above. In Table 1, we summarize their dependence on the system size $M$. In all cases, we can check that the product of the dimension of the subspaces $\mathcal{H}_{F}$, $\mathcal{H}_{F^{\prime}}$, and $\mathcal{H}_{\widetilde{D}}$ is equal to the dimension of the full Hilbert space, ${\rm dim}(\mathcal{H})=2^{N}$.
@@
$M$  $S=\lfloor\frac{M+2}{3}\rfloor$   $N_{C}$   $N_{P}$ $\leavevmode\nobreak\ S^{\prime}=\lfloor\frac{M+2}{6}\rfloor$
$6k$ $2k$ $1$ $13k-$ 1 $k$
$16k+$ 1 $12k+$ 1 00 $3k$ $k$
$26k+$ 2 $12k+$ 1 $1$ $3k$ $k$
$36k+$ 3 $12k+$ 1 $1$ $13k+$ 1 $k$
$46k+$ 4 $22k+$ 2 00 $13k+$ 1 $k+1$
$56k+$ 5 $22k+$ 2 $1$ $13k+$ 1 $k+1$
@@
IV Representation-dependent degeneracies: the subspace $\mathcal{H}_{\widetilde{D}}$
@@
While $[h_{m},\widetilde{h}_{n}]=0$ for all $m,n\in\{1,\ldots M\}$, only the first $M-1$ operators $\{\widetilde{h}_{m}\}_{m=1}^{M-1}$ commute with the boundary mode $\chi$, and hence with the physical fermion creation/annihilation operators. These $M-1$ operators generate the symmetry algebra $\mathcal{A}_{\widetilde{D}}$, which is the subject of this section.
In order to characterize $\mathcal{A}_{\widetilde{D}}$, we first observe that one can get a set of mutually commuting operators out of $\{\widetilde{h}_{m}\}_{m=1}^{M-1}$ by selecting one out of every three of them, say $\{\widetilde{h}_{1},\widetilde{h}_{4},\ldots\}$. This set is not yet maximal, as one can still add every other pair of consecutive operators out of the remaining $\widetilde{h}_{m}$, say $\{i\widetilde{h}_{2}\widetilde{h}_{3},i\widetilde{h}_{8}\widetilde{h}_{9},\ldots\}$. In total, this results in a set of $\sim M/3+M/6=M/2$ mutually commuting operators, whose eigenvalues ±$\pm 1$ account for a $2^{\sim M/2}$-fold degeneracy of the Hamiltonian spectrum.
@@
$[\tilde{X}_{n},\tilde{X}_{m}]=[\tilde{Y}_{n},\tilde{Y}_{m}]=0\,,\qquad[\tilde{X}_{n},\tilde{Y}_{m}]=2\delta_{n,m}\,,$ (21)
each pair $(\tilde{X}_{n},\tilde{Y}_{n})$ accounting for a two-fold degeneracy of the Hamiltonian spectrum. For some values of $M$ modulo $6$, these two families are completed by either one or two additional central operators, commuting with all $\tilde{X}_{m}$ and $\tilde{Y}_{m}$. Of these central operators, some belong exclusively to the algebra of the $\widetilde{h}_{m}$, while others can also be formulated as elements of the algebra of the $h_{m}$ (see paragraphs IV.5 and IV.6): for reasons that will be made clear below, the index $N_{C}$ in Table 1 counts only the former.
IV.1 $M=6k$
@@
The $13k-$ 1 operators $\tilde{X}_{n}$, the $13k-$ 1 operators $\tilde{Y}_{n}$ and the central operator $C$ generate the full algebra of the $\{\widetilde{h}_{m}\}_{m=1}^{M-1}$. They act on a Hilbert space of dimension $2^{3k}$, a basis of which can be defined by specifying the eigenvalues ±$\pm 1$ of all $\{\tilde{X}_{n}\}\cup C$, while the $\tilde{Y}_{n}$ flip those eigenvalues individually. In summary, the corresponding symmetry space has the structure
@@ IV.2 $M=6k+1$
while there is no central operator. As in the previous case, we can construct a family $\{\tilde{Y}_{n}\}$ obeying (21). These $M-1$ operators generate the full algebra of the $\{\widetilde{h}_{m}\}_{m=1\ldots M-1}$, and define a symmetry space of the form
@@
IV.3 $M=6k+2$
[…]
@@ IV.3 $M=6k+2$
In total, we then find $12\times 3k+1=M-$ 1 operators, generating the full algebra of the $\{\widetilde{h}_{m}\}_{m=1\ldots M-1}$. As in the previous section, those span a symmetry space with structure
@@ IV.4 $M=6k+3$
to which we can as usual associate the operators $\{\tilde{Y}_{n}\}$ satisfying (21). These $12\times(3k+1)=M-$ 1 generate the full algebra of the $\{\widetilde{h}_{m}\}_{m=1}^{M-1}$, and define a symmetry space of the form
@@
IV.5 $M=6k+4$
[…]
@@ IV.5 $M=6k+4$
Again, we can construct a family $\{\tilde{Y}_{n}\}$, obeying (21). We obtain a total of $12\times(3k+1)+1=M-$ 1 operators that generate the full algebra of the $\{\widetilde{h}_{m}\}_{m=1\ldots M-1}$. In contrast to the previously encountered cases where the central operator splits the Hilbert space $\mathcal{H}_{D}$ according to its ±$\pm 1$ eigenvalue, in the present case we simply write the corresponding symmetry space as

### v2

@@ 1 Introduction
Given the long history of the JW transformation, it is perhaps surprising that only recently new types of spin chains that can be mapped to free fermions were discovered. In 2019, Paul Fendley constructed a model which is quartic in the JW fermions, and still quadratic in terms of suitably defined fermionic creation and annihilation operators [19] (see Refs. [20, 21, 22] for earlier work in this direction). Remarkably, it was proven that the spin chain found by Fendley cannot be diagonalized by any generalized JW transformation [23], making the model fundamentally different from the previously known ones. This model and later generalizations [24, 25, 26, 27, 28] are now identified with the name of free fermions in disguise (FFD).
The mapping to free fermions in FFD models gives access to the full Hamiltonian spectrum. Still, the computation of spin correlation functions is generally hard, because the dictionary between states in the spin and free-fermion Hilbert spaces is highly non-trivial. Indeed, on the one hand the FFD mapping is non-linear and non-local; on the other hand, each Hamiltonian eigenspace is known to be exponentially degenerate [19], so that fermionic Fock states cannot be uniquely associated with spin states. As a consequence, the problem of computing correlation functions in FFD models remains essentially open, despite some recent progress in addressing the “inverse problem" of expressing local spin operators in terms of fermionic ones [29]. In some respect, the current situation is similar to that of interacting integrable models [30, 31]: while the Bethe ansatz gives direct access to the Hamiltonian spectrum, computing correlation functions in arbitrary eigenstates is hard, and analytic results exist only in a few special cases [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
@@
This work is organized as follows. We begin in Sec. 2 by recalling the model introduced by Paul Fendley (Sec. 2.1) and its exact solution in terms of free fermions (Sec. 2.2). Our results are contained in Secs. 2.3, 4, and 5, and summarized in Sec 3. In Sec. 2.3 we show that the Hilbert space factorizes as $\mathcal{H}=\mathcal{H}_{F}\otimes\mathcal{H}_{D}$, where $\mathcal{H}_{F}$ supports the fermionic operators. The structure of $\mathcal{H}_{D}$ is further resolved in the later sections, where we show that $\mathcal{H}_{D}=\mathcal{H}_{F^{\prime}}\otimes\mathcal{H}_{\widetilde{D}}$ and characterize the two factors individually. In particular, in Sec. 4 we show that $\mathcal{H}_{\widetilde{D}}$ is a “representation-dependent” subspace, generated by the common eigenstates of an extensive set of commuting Pauli strings, while in Sec. 5 we show that the space of operators supported on $\mathcal{H}_{F^{\prime}}$ is generated by an operator algebra of ancillary free fermion in disguise. Finally, Sec. 6 contains our conclusions and an outlook for future work.
2 The Model and its Hilbert-space structure
@@ 2.2 Exact solution
As an important property, transfer matrices at different spectral parameters commute, $T_{M}(u),T_{M}(u^{\prime})]=0$, and the Hamiltonian can obtained as the logarithmic derivative
@@
so that $T_{M}^{-1}(u)\propto T_{M}(-u)$. Here $P_{M}(u^{2})$ is a polynomial of degree $S=[(M+2)/3]$ in $u^{2}$ ($\cdot$] denotes the integer part), whose zeroes $\{u_{k}^{2}\}_{k=1,\ldots S}$ uniquely determine the energy spectrum of $H$: denoting $\epsilon_{k}=1/u_{k}$, there are $2^{S}$ distinct energy eigenvalues of the form
@@
$\Psi_{k}\propto T_{M}(\epsilon_{k})~\chi~T_{M}(-\epsilon_{k})\,.$ (7)
@@
2.3 The Hilbert-space structure
In this section, we use the fermionic operators (7) to show that the FFD Hilbert space can be decomposed as $\mathcal{H}=\mathcal{H}_{F}\otimes\mathcal{H}_{D}$. Here, $\mathcal{H}_{F}$ is a space of dimension $2^{S}$ that the fermionic operators are supported on, while $\mathcal{H}_{D}$ accounts for the exponentially large degeneracies observed in the spectrum of $H$. This simple observation is our first result and the basis for the subsequent analysis presented in the rest of this work, cf. Sec. 3
@@ 2.3 The Hilbert-space structure
[…]
@@
[…]
Next, we claim that the isomorphism $\varphi$ maps the operators $\Psi_{k}$ into operators supported on $\mathcal{H}_{F}$, namely $\varphi(\Psi_{k})=c_{k}\otimes\mathbb{1}$. Indeed, by definition of $\varphi$,
@@
so that all fermionic operators are supported on $\mathcal{H}_{F}$. Finally, since the Hamiltonian is written explicitly in terms of the fermionic operators, it follows that (with a slight abuse of notation) $H=H\otimes\mathbb{1}$. Therefore, $\mathcal{H}_{D}$ accounts for the exponential degeneracy of each Hamiltonian eigenstate.
3 Overview of our results
We have shown that the Hilbert space admits the factorization $\mathcal{H}=\mathcal{H}_{F}\otimes\mathcal{H}_{D}$. In the rest of this work, we will further refine the structure of $\mathcal{H}_{D}$. Before proceeding, it is useful to provide an overview of our main results, which we report in this section.
@@ 3 Overview of our results
  • We will show that ${\mathcal{H}}_{F^{\prime}}$ supports the action of all operators written in terms of the generators $h_{m}$ that commute with the fermion creation/annihilation operators. We will denote the algebra of such operators by $\mathcal{A}_{F^{\prime}}$ and we will refer to it as the ancillary fermionic algebra (to be distinguished by the fermionic algebra $\mathcal{A}_{F}$ supported on $\mathcal{H}_{F})$. Importantly, the definition of $\mathcal{A}_{F^{\prime}}$ is independent of the specific representation chosen, because it can be stated entirely in terms of the generators $h_{m}$ and their algebraic relations. However, we will show that $\mathcal{A}_{F^{\prime}}$ does depend on the Hamiltonian couplings $\{b_{m}\}$. Going further, we will show in Sec. 5, that $\mathcal{A}_{F^{\prime}}$ is a fermionic algebra, generated by a new family of ancillary fermion operators
@@
The dimensions of the subspaces $\mathcal{H}_{F}$, $\mathcal{H}_{F^{\prime}}$, and $\mathcal{H}_{\widetilde{D}}$ depend, respectively, on the integers $S$, $S^{\prime}$, $N_{P}$ and $N_{C}$ introduced above. In Table 1, we summarize their dependence on the system size $M$. In all cases (except for the case $M=6k+3$, where an additional zero mode has to be included to the construction, cf. the end of Sec. 5), we can check that the product of the dimension of the subspaces $\mathcal{H}_{F}$, $\mathcal{H}_{F^{\prime}}$, and $\mathcal{H}_{\widetilde{D}}$ is equal to the dimension of the full Hilbert space, ${\rm dim}(\mathcal{H})=2^{N}$.
@@
$M$  $S=\lfloor\frac{M+2}{3}\rfloor$   $N_{C}$   $N_{P}$ $S^{\prime}=\lfloor\frac{M+2}{6}\rfloor$
$6k$ $2k$ 11 3k13k-1 $k$
$6k+1$ $2k+1$ 0 $3k$ $k$
$6k+2$ $2k+1$ 11 $3k$ $k$
$6k+3$ $2k+1$ 0 $3k+1$ $k$
$6k+4$ $2k+2$ 0 $3k+1$ $k+1$
$6k+5$ $2k+2$ 11 $3k+1$ $k+1$
@@
4 Representation-dependent degeneracies: the subspace $\mathcal{H}_{\widetilde{D}}$
@@
While $h_{m},\widetilde{h}_{n}]=0$ for all $m,n\in\{1,\ldots M\}$, only the first M1M-1 operators $\{\widetilde{h}_{m}\}_{m=1}^{M-1}$ commute with the boundary mode $\chi$, and hence with the physical fermion creation/annihilation operators. These M1M-1 operators generate the symmetry algebra $\mathcal{A}_{\widetilde{D}}$, which is the subject of this section.
In order to characterize $\mathcal{A}_{\widetilde{D}}$, we first observe that one can get a set of mutually commuting operators out of $\{\widetilde{h}_{m}\}_{m=1}^{M-1}$ by selecting one out of every three of them, say $\{\widetilde{h}_{1},\widetilde{h}_{4},\ldots\}$. This set is not yet maximal, as one can still add every other pair of consecutive operators out of the remaining $\widetilde{h}_{m}$, say $\{i\widetilde{h}_{2}\widetilde{h}_{3},i\widetilde{h}_{8}\widetilde{h}_{9},\ldots\}$. In total, this results in a set of $\sim M/3+M/6=M/2$ mutually commuting operators, whose eigenvalues $\pm 1$ account for a $2^{\sim M/2}$-fold degeneracy of the Hamiltonian spectrum.
@@
$\tilde{X}_{n},\tilde{X}_{m}]=[\tilde{Y}_{n},\tilde{Y}_{m}]=0\,,\qquad[\tilde{X}_{n},\tilde{Y}_{m}]=2\delta_{n,m}\,$, (21)
each pair $(\tilde{X}_{n},\tilde{Y}_{n})$ accounting for a two-fold degeneracy of the Hamiltonian spectrum. For some values of $M$ modulo $6$, these two families are completed by either one or two additional central operators, commuting with all $\tilde{X}_{m}$ and $\tilde{Y}_{m}$. Of these central operators, some belong exclusively to the algebra of the $\widetilde{h}_{m}$, while others can also be formulated as elements of the algebra of the $h_{m}$ (see paragraphs 4.5 and 4.6): for reasons that will be made clear below, the index $N_{C}$ in Table 1 counts only the former.
4.1 $M=6k$
@@
The 3k13k-1 operators $\tilde{X}_{n}$, the 3k13k-1 operators $\tilde{Y}_{n}$ and the central operator $C$ generate the full algebra of the $\{\widetilde{h}_{m}\}_{m=1}^{M-1}$. They act on a Hilbert space of dimension $2^{3k}$, a basis of which can be defined by specifying the eigenvalues $\pm 1$ of all $\{\tilde{X}_{n}\}\cup C$, while the $\tilde{Y}_{n}$ flip those eigenvalues individually. In summary, the corresponding symmetry space has the structure
@@ 4.2 $M=6k+1$
while there is no central operator. As in the previous case, we can construct a family $\{\tilde{Y}_{n}\}$ obeying (21). These M1M-1 operators generate the full algebra of the $\{\widetilde{h}_{m}\}_{m=1\ldots M-1}$, and define a symmetry space of the form
@@
4.3 $M=6k+2$
When $M\ {\rm mod}\ 6=2$, a maximal set of mutually commuting operators is given by
@@ 4.3 $M=6k+2$
In total, we then find $2\times 3k+1=M-1$ operators, generating the full algebra of the $\{\widetilde{h}_{m}\}_{m=1\ldots M-1}$. As in the previous section, those span a symmetry space with structure
@@ 4.4 $M=6k+3$
to which we can as usual associate the operators $\{\tilde{Y}_{n}\}$ satisfying (21). These $2\times(3k+1)=M-1$ generate the full algebra of the $\{\widetilde{h}_{m}\}_{m=1}^{M-1}$, and define a symmetry space of the form
@@
4.5 $M=6k+4$
When $M\ {\rm mod}\ 6=4$, a maximal set of mutually commuting operators can be constructed as
@@ 4.5 $M=6k+4$
Again, we can construct a family $\{\tilde{Y}_{n}\}$, obeying (21). We obtain a total of $2\times(3k+1)+1=M-1$ operators that generate the full algebra of the $\{\widetilde{h}_{m}\}_{m=1\ldots M-1}$. In contrast to the previously encountered cases where the central operator splits the Hilbert space $\mathcal{H}_{D}$ according to its $\pm 1$ eigenvalue, in the present case we simply write the corresponding symmetry space as